OpenStudy (anonymous):

a student has test scores of 67,76, and 81 points. What must she score on the fourth exam to have an average score of at least 75 points? Express the solution as an inequality. Use s as the unknown variable.

OpenStudy (anonymous):

(67+76+81+s) / 4 = 75 (224+s) / 4 = 75 75 * 4 = 300 224+s = 300 300-224=76 s = 76

OpenStudy (anonymous):

Omgggg thankyou!!! :D